Rocket Fuel in Drinking Water. Near Las Vegas, NV, improper disposal of perchlorates used to manufacture rocket fuel has contaminated a stream that flows into Lake Mead, the largest artifical lake in the United States and a major supply of drinking and irrigation water for the American Southwest. The EPA has proposed an advisory range for perchlorate concentrations in drinking water of 4 to 18 micrograms/Liter. the perchlorate concentration in the stream averages 700 micrograms/ liter, and the stream flows at an average rate of 161 million gallons per day (1 gallon=3.785L) a. What are the formlas of sodium perchlorate and ammonium perchlorate? b. How many kilograms of perchlorate flow from the Las Vegas stream into Lake Mead each day? c. What volume of perchlorate - free lake water would have to mix with the stream water each dayto dilute the stream's perchlorate concentration from 700 to 4 micrograms/Liter
Answer #1

a. What are the formlas of sodium perchlorate and ammonium perchlorate?

Sodium ion: Na+

perchlorate ion: ClO4-

sodium perchlorate: NaClO4

Ammonium ion: NH4+

Ammonium perchlorate : NH4ClO4

b. How many kilograms of perchlorate flow from the Las Vegas stream into Lake Mead each day?

stream flows at an average rate = 161 million gallons/day (1 gallon=3.785L)

Average perchlorate concentration in the stream =700 \mu g/L

Mass of perchlorate flow into Lake mead =

\frac{161\times 10^6 gallons}{day}\times \frac{3.785 L}{1 gallon} = 609.4\times 10^6 L/day

\frac{609.4\times 10^6 L}{day}\times \frac{700\times 10^{-6}g}{1L}\times \frac{1kg}{1000g}

=426.6 kg/day

c. What volume of perchlorate - free lake water would have to mix with the stream water each dayto dilute the stream's perchlorate concentration from 700 to 4 micrograms/Liter

Formula

C1V1 = C2V2

COncnetration of perchlorate inn stream = M1= 700 x 10-6 g/L

Volume of stream flow =  609.4 x 106 L

Concentration of lake should be = 4 micrograms/Liter =  4 x 10-6 g/L

700 x 10-6 g/L x 609.4 x 106 L = 4 x 10-6 g/L x V2

V_2 = \frac{426569.5}{4\times 10^{-6}} L

V2 = 1.066 x 1011 L

Volume of perchlorate free water have to be mixed =

1.066 x 1011 L - 609.4 x 106 L

= 1.06 x 1011 L water has to be mixed with stream